OpenStudy (anonymous):

On Monday morning, Colin filled his car with six and one-half gallons of gas, then used three and one-fourth gallons driving. On Tuesday, he used up 2 gallons driving. On Tuesday night, his gas tank had four and one-fourth gallons remaining. How much gas was in Colin's tank before he filled up on Monday morning?

OpenStudy (allank):

Let the amount of gas in the tank before he filled up on Monday be x gallons. Activity : Amount of gas in tank Monday morning: x+6.5 Monday driving: x+6.5-3.25=x+3.25 Tuesday driving: x+3.25-2= x+1.25 = Final amount of gas. Now, use this sentence to form an equation: "On Tuesday night, his gas tank had four and one-fourth gallons remaining"